# Preference Robustness for DPO with Applications to Public Health

### v1

@@ 5.1 Benchmarks on General Alignment Datasets
We use the recent UltraFeedback alignment benchmark dataset (Cui et al. 2024), consisting of 60,000 top-quality human pairwise preference annotations that cover a wide variety of tasks and instruction types, rendering it a broad benchmark for examining alignment research.
@@
Our primary objective is to evaluate how effectively preference-based LLM fine-tuning methods handle shifts in preference distributions between training and evaluation data. We use $q^{*}$ denote the true preference distribution used at test time. For our experiments, we define $q^{*}$ using the state-of-the-art preference reward model Eurus 7B(Yuan et al. 2025), which has been trained for the UltraFeedback dataset (Cui et al. 2024). We compute $q^{*}$ using the BT model.
@@
Intuitively, $q_{\alpha}$ “flips” the preference distribution with probability $\alpha$. We refer to this type of noise interchangeably as label switching. During training, we have access to $q_{\alpha}$ but not to $q^{*}$, with $\alpha$ being unknown. We include additional experiments with alternative forms of noise in Appendix.
@@
We base our experiments on the recent Phi 3-mini 3B model without instruction (Abdin et al. 2024). We initially train for two epochs of supervised fine tuning (SFT), followed by one epoch of DPO, with the baselines discussed below. All experiments are performed on a Nvidia A100 GPU 40 GB with batch size 2. Refer to the technical appendix for full details on the training procedure such as weight decay, learning rate scheduler, low-rank adaptation parameters, quantization, etc and experiments conducted on the Llama 3 8B model (Grattafiori et al. 2024), not shown here due to space limitations.
@@
We compare DPO-PRO against vanilla DPO (Rafailov et al. 2023) and the two robust baselines: rDPO (Chowdhury, Kini, and Natarajan 2024) and DrDPO (Wu et al. 2025). We implement DPO-PRO with a value of $\rho=0.1$ for the $\chi^{2}$-divergence. We train for two noise levels ($\alpha=0.1$ and $\alpha=0.3$), termed low and high respectively, and also for the noiseless case $\alpha=0$.
Evaluation metrics
After the DPO step, we generate a response for every prompt $x$ in the evaluation dataset $y_{g}\sim\pi_{\theta}(x)$. Then, we compute the reward $R_{\phi}(x,y_{g})$. We focus on two evaluation metrics, the win rate, which is computed by comparing the generated prompt with the chosen response $y_{c}$ in the dataset as $\texttt{win\_rate}=\frac{1}{N_{\mathrm{eval}}}\sum_{i=1}^{N_{\mathrm{eval}}}\mathbb{I}(R_{\phi}(x,y_{g})>R_{\phi}(x,y_{c})),$ where $\mathbb{I}$ is the indicator function. The next metric is the generation evaluation reward, computed simply as $\texttt{eval\_reward}=\frac{1}{N_{\mathrm{eval}}}\sum_{i=1}^{N_{\mathrm{eval}}}R_{\phi}(x,y_{g})$.
@@ Results
Table 1(b) presents the results. Across all noise levels, DPO-PRO consistently achieves highest performance in both win rate and evaluation reward. Notably, it outperforms all baselines even in noise-free setting, highlighting robustness without sacrificing standard performance.
We found it surprising that performance of all methods increased with additional noise. This can be explained by the fact that DPO in fixed (off-policy) data is known to be highly sensitive to reward hacking (Gheshlaghi Azar et al. 2024; Amini, Vieira, and Cotterell 2024; Yang et al. 2024; Xiao et al. 2025). An analogous result for DPO-based methods can be found in Wu et al. (2025); Bukharin et al. (2024). The noise seemingly acts as a regularization, preventing model from overfitting to spurious patterns in training data and instead learning more generalizable features. This phenomenon is consistent with recent theoretical work showing controlled noise injection can improve generalization by encouraging simpler decision boundaries (Li et al. 2025).
Table 1: Results on UltraFeedback under label-flip noise. Experiments run with Phi-3-mini. Mean and standard error over 2000 evaluation samples are shown.
(a) Win rate ($\uparrow$)
Method No noise Low noise (0.1) High noise (0.3)
[…]
(b) Generation Evaluation Reward ($\uparrow$)
Method No noise Low noise (0.1) High noise (0.3)
[…]
@@ 5.2 Benchmarks on Maternal Health Data
We apply our proposed method in the context of mobile health program run by ARMMAN (2025). Specifically, we have anonymized beneficiary listenership data from a quality-improvement study conducted by ARMMAN [^1] The authors have been access to this restricted dataset under a data usage agreement.. The data consists of registration information of beneficiaries (mothers) containing their socio-demographic features of age, income education as well as their preferences on call slot times and language. Additionally, we have information on beneficiaries’ interaction with automated voice calls and live service calls made by health-workers. We model the problem of allocating live service calls as a RMAB problem. For all our experiments, we consider a sample of population with $2100$ beneficiaries and $K=210$ as the budget, representing the number of live service calls that can be made every week. Exact policy optimization for RMABs is generally known to be intractable (Papadimitriou and Tsitsiklis 1987). However, Whittle index policy (Whittle 1988) is a popular heuristic to obtain an asymptotically optimal policy. Please refer also to the technical appendix for additional details.
@@
We generate a preference dataset of health worker prioritization commands, preferred, and rejected reward functions as shown in Figure 1. Since it is costly to generate this dataset through human annotation, we use an LLM judge with ChatGPT 4o-mini (OpenAI 2024) as follows: i) query LLM-judge to obtain 20 candidate reward functions that align with the prioritization command, ii) for each candidate reward function, solve RMAB problem using Whittle index method and generate trajectory outputs, iii) sample 50 pairs of reward functions from this set and query the LLM-judge to select preferred reward function, iv) perform this query 10 times to estimate the LLM’s uncertainty over preferences. Finally, we obtain a dataset of 9500 preferred and rejected reward function responses over 190 prioritization commands (see Appendix for more details).
Noise injection, training details, and baselines
We adopt the same noise injection scheme to obtain $q_{\alpha}$ as described in Section 5.1 and Eq (4). We fine-tune a Llama 3 (8B) model, keeping the learning rate, DPO reward temperature, $\chi^{2}$-divergence radius, and other hyperparameters identical to those used in the UltraFeedback benchmark experiments. Given poor performance of rDPO in Table 1(b), we omit this baseline in this experiment.
@@
To evaluate the quality of a generated reward function, we follow a procedure analogous to the dataset creation: i) solve the resulting RMAB problem using the Whittle Index method to generate policy trajectories, ii) use LLM judge (GPT-4o-mini) to evaluate these policy outcomes. Specifically, we compute the win rate by comparing the policy induced by our generated reward function to the policy obtained from the reward function in the training data.
@@
Figure 3: Winrates for RMAB environment under label flipping noise. We present more results in the Appendix.
Results
Table 3 shows the results. In zero-noise setting, vanilla-DPO outperforms robust methods. However, as noise level increases, DPO-PRO outperforms vanilla DPO and DrDPO. We also perform a task-level analysis of winrates in Figure 3. We note that across all tasks, under the presence of noise, robust DRO methods are better than vanilla-DPO. Especially, in 5 out of 8 evaluation task categories, DPO-PRO outperforms baselines under the presence of noise. This signifies the robustness of our approach.
To further illustrate the differences between vanilla DPO and DPO-PRO, we present response outputs from both models on two representative tasks from the evaluation dataset (Table 2). Many tasks contain attributes that can be interpreted in multiple ways. For instance, the concept of “young beneficiaries” can be identified directly through the age feature, but also indirectly via lower education level or income. Similarly, a task that expresses a preference for “midday calls” could refer explicitly to the 12:30–3:00 PM time slot, but might also include the 10:30–12:30 PM window, depending on interpretation.
5.3 Comparison with Decision-Language Model
Table 4 compares DPO-PRO with the Decision Language Model (DLM), a self-reflection-based approach introduced in (Behari et al. 2024), using two evaluation metrics: (1) win rate, determined by an external LLM judge (GPT-4o-mini), and (2) reward, computed using ground-truth human-designed reward functions. While DLM slightly outperforms DPO-PRO on both metrics, the results indicate that DPO-PRO achieves comparable alignment quality despite its simplicity and significantly lower inference-time cost.
Figure 2 provides detailed comparison of inference-time cost between DPO-PRO and DLM. Unlike DLM, whose inference cost scales linearly with number of arms due to repeated simulation and self-reflection, DPO-PRO incurs a constant inference cost regardless of problem size. This efficiency gap becomes increasingly significant in large-scale settings, making DPO-PRO more scalable. While overall performance is comparable or mixed across evaluation metrics, minimal inference-time overhead makes DPO-PRO especially well-suited for real-world deployment.
Table 3: Winrate (%, $\uparrow$) comparision for ARMMAN environment for different methods under no noise and high noise settings, aggregated across all tasks.
Method Vanilla-DPO DrDPO DPO-PRO
No Noise $\bf 55\pm 2.3$ $40\pm 1.85$ $35.3\pm 1.83$
High Noise (0.3) $45\pm 2.03$ $54.8\pm 1.84$ $\bf 63.9\pm 1.94$
Table 4: Comparison of Winrate and Reward
Method Winrate Ground Truth Reward
DLM $35.5\pm 5.02$ $6073.04\pm 88.5$
DPO-PRO $32.2\pm 4.92$ $5687.3\pm 112.3$

### v2

@@ 5.1 Benchmarks on General Alignment Datasets
We use the recent UltraFeedback alignment benchmark dataset (ultra2024), consisting of 60,000 top-quality human pairwise preference annotations that cover a wide variety of tasks and instruction types, rendering it a broad benchmark for examining alignment research.
@@
Our primary objective is to evaluate how effectively preference-based LLM fine-tuning methods handle shifts in preference distributions between training and evaluation data. We use $q^{*}$ denote the true preference distribution used at test time. For our experiments, we define $q^{*}$ using the state-of-the-art preference reward model Eurus 7B(yuan2025eurus), which has been trained for the UltraFeedback dataset (ultra2024). We compute $q^{*}$ using the BT model.
@@
Intuitively, $q_{\alpha}$ “flips” the preference distribution with probability $\alpha$. We refer to this type of noise interchangeably as label switching. During training, we have access to $q_{\alpha}$ but not to $q^{*}$, with $\alpha$ being unknown.
@@
We base our experiments on the recent Phi 3-mini 3B model without instruction (abdin2024phi3). We initially train for two epochs of supervised fine tuning (SFT), followed by one epoch of DPO, with the baselines discussed below. All experiments are performed on a Nvidia A100 GPU 40 GB with batch size 2.
@@
We compare DPO-PRO against the standard DPO (rafailov2023DPO) and the prior distributionally robust variant DrDPO (wu2025drdpo) using their best hyperparameters suggested by the authors. We evaluate DPO-PRO under several values of the robustness parameter $\rho$ based on the $\chi^{2}$-divergence. Models are trained under two levels of label-flip noise, $\alpha=0.3$ (low) and $\alpha=0.6$ (high), as well as the noiseless setting $\alpha=0$.
Evaluation
After the fine-tuning step, we generate a response $y_{g}\sim\pi_{\theta}(x)$ for every prompt $x$ in the evaluation dataset and compute its reward $R_{\phi}(x,y_{g})$. We consider two evaluation metrics. The first is the win rate, computed by comparing the generated response with the chosen response $y_{c}$ in the dataset using an LLM-as-a-judge approach with GPT-4o-mini (openai_gpt4omini_2024). To mitigate position bias, we randomly flip the order of the two responses presented to the judge. The second metric, evaluation reward, is computed as the average reward score defined as $\frac{1}{N_{\mathrm{eval}}}\sum_{i=1}^{N_{\mathrm{eval}}}R_{\phi}(x,y_{g}).$
@@ Results
Table 1 summarizes the results. For the LLM-as-a-judge evaluation, the different DPO-PRO variants (each corresponding to different $\rho$ value) tend to underperform the baselines in the low-noise setting, but their relative performance improves as noise increases. Under high noise, all DPO-PRO variants outperform both vanilla DPO and DrDPO. The effect of $\rho$ follows the expected pattern: smaller $\rho$ values perform well when noise is low but degrade more quickly as noise increases, while larger $\rho$ values exhibit more stability. Overall, the DPO-PRO family shows improved robustness as the preference noise becomes larger.
For metrics based on the reward model, DPO-PRO outperforms the baselines even in the noiseless setting, except for the case with $\rho=0.1$. As expected, performance declines for all methods as noise increases. In the low-noise regime, however, the DPO-PRO variants perform slightly worse than the baseline, but they regain and ultimately surpass baseline performance as the noise becomes even larger. Note that this does not contradict the DRO objective, because the true optimization target for fine-tuning is reward plus KL regularization, while our evaluation here relies solely on reward-model scores. Moreover, these scores are only a proxy for quality, particularly for out-of-distribution responses, and are limited by the reward model’s own generalization capabilities.
(a) Win rate (%) based on LLM-as-a-judge evaluation ($\uparrow$).
Method No noise Low noise High noise
DPO-PRO ($\rho=0.008$) 43.24 35.17 28.75
DPO-PRO ($\rho=0.03$) 41.71 33.33 28.98
DPO-PRO ($\rho=0.1$) 40.33 33.54 27.97
DPO 43.77 36.19 26.97
DrDPO 44.85 38.84 25.51
(b) Evaluation reward ($\uparrow$).
Method No noise Low noise High noise
DPO-PRO ($\rho=0.008$) 923.70 713.00 446.65
DPO-PRO ($\rho=0.03$) 912.65 695.00 499.59
DPO-PRO ($\rho=0.1$) 781.28 656.32 499.36
DPO 877.01 729.12 400.44
DrDPO 897.67 753.77 419.05
Table 1: Results on UltraFeedback under label-flip noise.
@@ 5.2 Benchmarks on Maternal Health Data
We apply our proposed method in the context of mobile health program run by ARMMAN. We use real-world anonymized beneficiary listenership data from a quality-improvement study conducted by ARMMAN [^1] The authors have been access to this restricted dataset under a data usage agreement.. The data consists of registration information of beneficiaries (mothers) containing their socio-demographic features of age, income education as well as their preferences on call slot times and language. Additionally, we have information on beneficiaries’ interaction with automated voice calls and live service calls made by health-workers. We model the problem of allocating live service calls as a RMAB problem. For all our experiments, we consider a population with $2100$ beneficiaries (i.e., 2100 arms) and the number of live service calls that can be made every week is limited to 210.
We use the Phi 3-Mini 3B model, and the noise-injection and training procedures follow the same setup as in Section 5.1. Here, we experiment only with the noiseless and low-noise settings to better reflect realistic conditions, rather than injecting artificially large noise.
@@
We generate a preference dataset of health worker prioritization commands, preferred, and rejected reward functions. Since it is costly to generate this dataset through human annotation, we use an LLM judge with ChatGPT 4o-mini (openai_gpt4omini_2024) as follows: i) query LLM-judge to obtain 20 candidate reward functions that align with the prioritization command, ii) for each candidate reward function, solve RMAB problem using Whittle index method (whittle1988restless) and generate trajectory outputs (see (math11071639) or the Appendix in (kim2025dpopro) for the explanation on the Whittle index), iii) sample 50 pairs of reward functions from this set and query the LLM-judge to select preferred reward function, iv) perform this query 10 times to estimate the LLM’s uncertainty over preferences. Finally, we obtain a dataset of 9500 preferred and rejected reward function responses over 190 prioritization prompts.
Baselines
In addition to the DPO variants, we also compare against the Decision Language Model (DLM), a self-reflection–based inference-time approach introduced in (behari2024a) using Gemini 2.0. Note that DLM is not fine-tuned on the training data, so this method has no notion of noise. For each prompt, we simply apply the out-of-the-box Gemini model and run the self-reflection procedure described in (behari2024a). We use Gemini 2.0, as in the original work, rather than the open-source Phi-3B model used for fine-tuning. This is to assess how our method compares against a state-of-the-art closed model that a typical user would rely on, further strengthened by self-reflection.
@@
To evaluate the quality of each generated reward function, we follow a procedure parallel to the dataset construction. For every prompt in the evaluation set, we generate a reward function using a given fine-tuning method and obtain the corresponding RMAB policy using the Whittle Index. We repeat the same procedure using the chosen reward function associated with that prompt in the evaluation data, producing a reference policy for comparison.
We compare these policies using two complementary methods. First, because this task provides access to ground-truth human-designed reward functions from behari2024a, we can directly evaluate RMAB policy quality by computing the achieved reward under this ground-truth function. We compare the average ground-truth reward obtained by the policy induced by the fine-tuned LLM with that obtained by the policy from the evaluation data. Second, we also perform an LLM-as-a-judge evaluation using GPT-4o-mini to compute win-rate as in Section 5.1. The judge compares the resulting trajectories of the RMAB policies, i.e., the resource allocation outcomes across the demographic data, to determine which policy is preferred.
Results
We report the results in Table 2. For the ground-truth reward metric, all variants of DPO-PRO outperform both DrDPO and Vanilla DPO in both the noiseless and high-noise settings. Their performance is slightly lower than DLM, but this gap must be interpreted in context: DLM requires (1) access to a high-end proprietary LLM API at deployment time and (2) substantially higher inference-time cost, as we show below.
For the win-rate metric, the results are more mixed. The DPO-PRO variants generally outperform DrDPO and Vanilla DPO in the low-noise setting, with the exception of the $\rho=0.03$ variant. In particular, DPO-PRO ($\rho=0.1$) outperforms both DrDPO and Vanilla DPO in both the noiseless and low-noise settings, while DPO-PRO ($\rho=0.3$) performs similarly to these baselines under no noise and clearly exceeds them under low noise. As with the ground-truth reward metric, all DPO-PRO variants perform below DLM.
One interesting observation is that, unlike the clean trends in Section 5.1, the performance across different $\rho$ values and noise levels in the ARMMAN task does not yield a simple or monotonic pattern. For example, larger $\rho$ values are not consistently more conservative, and performance even improves as noise is injected. Although pinpointing the exact cause is difficult, several factors likely contribute to this behavior.
First, unlike the curated benchmark data in Section 5.1, the ARMMAN dataset is substantially noisier. Preference labels generated by LLM judges are difficult to obtain reliably for this task (see Section 3), reducing the quality of the signal available during fine-tuning. In other words, the noiseless setting here does not necessarily correspond to a clean preference signal, since the underlying annotations are already highly noisy. Second, DPO and related off-policy preference-based methods are known to be sensitive to overfitting and reward hacking when trained on fixed datasets (pmlr-v238-gheshlaghi-azar24a; amini-etal-2024-direct; yang2024orthogonalfinetuningdirectpreference; xiao2025comprehensivesurveydirectpreference). In such settings, injected noise can act as a regularizer: it discourages the model from fitting spurious or inconsistent preference patterns and promotes more stable decision rules. Similar trends, where performance improves under moderate noise, have been reported in (wu2025drdpo; bukharin2024robust). This interpretation is also consistent with recent theoretical work showing that controlled noise injection can improve generalization by implicitly encouraging simpler and more robust decision boundaries (li2025towards).
Inference Time Comparison
Figure 1 compares the inference time of DPO-PRO (with $\rho=0.1$) and DLM. Unlike DLM, whose inference time scales linearly with the number of arms due to repeated simulation of resource-allocation policy and self-reflection, DPO-PRO incurs a constant inference cost regardless of problem size. This efficiency gap becomes increasingly significant in large-scale settings, making DPO-PRO more scalable. While overall performance is comparable or mixed across evaluation metrics, minimal inference-time overhead makes DPO-PRO especially well-suited for real-world deployment.
Examples
To further illustrate the differences between vanilla DPO and DPO-PRO (with $\rho=0.1)$, we present response outputs from both models on two representative tasks from the evaluation dataset (Table 3). Many tasks contain attributes that can be interpreted in multiple ways. For instance, the concept of “young beneficiaries” can be identified directly through the age feature, but also indirectly via lower education level or income. Similarly, a task that expresses a preference for “midday calls” could refer explicitly to the 12:30–3:00 PM time slot, but might also include the 10:30–12:30 PM window, depending on interpretation.
(a) Average Ground Truth Reward ($\uparrow$).
Method No noise Low noise
DLM $6073.04$ NA
DPO $5465.36$ $5707.09$
DrDPO $5480.06$ $5675.17$
DPO-PRO ($\rho=0.008$) $5530.58$ $5832.66$
DPO-PRO ($\rho=0.03$) $5557.9$ $5790.8$
DPO-PRO ($\rho=0.1$) $5677.76$ $5807.83$
DPO-PRO ($\rho=0.3$) $5797.1$ $5904.7$
(b) Win rate (%, $\uparrow$).
Method No noise Low noise
DLM 24.82 NA
DPO 15.38 13.84
DrDPO 15.89 12.13
DPO-PRO ($\rho=0.008$) 14.18 14.87
DPO-PRO ($\rho=0.03$) 13.5 13.5
DPO-PRO ($\rho=0.1$) 17.7 14.35
DPO-PRO ($\rho=0.3$) 15.7 17.8
Table 2: Results on ARMMAN environment under no-noise and low-noise settings.
Figure 1: Inference-time comparison of DPO-PRO and DLM across different population sizes (i.e., number of arms) in the underlying RMAB.